### v1

# ViTSGMM: A Robust Semi-Supervised Image Recognition Network Using Sparse Labels

Abstract
We present ViTSGMM, an image recognition network that leverages semi-supervised learning in a highly efficient manner. Existing works often rely on complex training techniques and architectures, while their generalization ability when dealing with extremely limited labeled data remains to be improved. To address these limitations, we construct a hierarchical mixture density classification decision mechanism by optimizing mutual information between feature representations and target classes, compressing redundant information while retaining crucial discriminative components. Experimental results demonstrate that our method achieves state-of-the-art performance on STL-10 and CIFAR-10/100 datasets when using negligible labeled samples. Notably, this paper also reveals a long-overlooked data leakage issue in the STL-10 dataset for semi-supervised learning tasks and removes duplicates to ensure the reliability of experimental results.  [^1] Code available at https://github.com/Shu1L0n9/ViTSGMM.
Keywords: Semi-Supervised Learning, Image Recognition, Vision Transformer, Gaussian Mixture Model
1 Introduction
Deep learning has achieved remarkable success in image classification tasks. However, its performance is highly dependent on large-scale annotated datasets. In real-world applications, labeled data is often scarce and expensive to acquire. Therefore, semi-supervised learning (SSL) has received widespread attention in recent years as an effective solution. The goal of semi-supervised learning is to train models using a small amount of labeled data and large amount of unlabeled data, thereby improving the model’s generalization ability. Nevertheless, most existing semi-supervised learning methods, such as consistency regularization and pseudo-labeling, rely on complex training techniques and architectural designs, and their generalization ability when the labeled datas are extremely limited remains to be improved. Furthermore, although some studies have attempted to apply the recently popular Vision Transformer (ViT) to semi-supervised learning, they have not yet fully leveraged its advantages.
Gaussian Mixture Models (GMMs) are powerful probabilistic models that assume data is generated from a mixture of multiple Gaussian distributions. In image recognition tasks, GMMs can model the distribution of data features and subsequently perform classification. The inspiration for this research comes from Miller & Uyar’s systematic study of the “Mixture of Experts” structure based on the joint probability model of features and labels. In our research, we found that the performance of GMMs is highly related to the quality of feature vectors. If the features lack representativeness or have insufficient discriminative power, the classification performance will be severely affected.
To address this, we designed a feature encoding framework based on a self-attention mechanism to fully exploit the global contextual information in images and generate more discriminative feature representations. Simultaneously, we incorporated semi-supervised learning strategies, utilizing a pseudo-labeling mechanism to efficiently leverage unlabeled data. This approach achieves superior classification performance within the framework of mixture density estimation. This method significantly overcomes the limitations of traditional methods in terms of feature representation capability and model generalization performance.
Taking the Dino with ViT-large backbone as our feature extraction model, we validated the proposed semi-supervised Gaussian mixture model (SGMM) method on three commonly used image classification datasets: STL-10, CIFAR-10, and CIFAR-100. The experimental results show that our method achieves state-of-the-art performance and exhibits good stability and robustness.
Figure 1: Performance Comparison. Comparison of our method with recent semi-supervised learning models on STL-10 dataset with different numbers of labels, achieving state-of-the-art performance.
It is worth mentioning that we have revealed a serious problem that has been overlooked by existing semi-supervised learning research: there are 7,500+ duplicate samples between the unlabeled samples of STL-10 and the test set samples, which is undoubtedly a serious data leakage problem when directly used for semi-supervised learning tasks.
Our main contributions can be summarized as follows:
  • We introduce the powerful Vision Transformer (ViT) as a feature extractor into the classical Mixture of Experts (MoE) classifier for image recognition tasks.
  • We design a simple but performant semi-supervised learning strategy that effectively leverages negligible labels.
  • We demonstrate that this method achieves state-of-the-art performance.
  • We reveal the longstanding, yet overlooked issue of trainset and testset contamination in the STL-10 dataset, which has not been addressed in other works.
The remainder of this paper is organized as follows: Section 2 reviews related works; Section 3 introduces our ViTSGMM network; Section 4 demonstrates and analyzes experimental results; Section 5 discusses the significance and limitations of the research findings and outlines future research directions.
2 Related Works
2.1 Semi-Supervised Learning and Gaussian Mixture Model (GMM)
Semi-Supervised Learning (SSL) has been widely applied in tasks such as image classification and speech recognition in recent years, especially in scenarios where labeled data is scarce. Early SSL methods mainly relied on generating pseudo-labels and consistency regularization.
Pseudo-Labeling is one of the classic methods in SSL, with the core idea of using an existing model to predict unlabeled data Yarowsky (1995) and adding its predicted labels to the training set as pseudo-labels. Representative works include Lee et al. (2013) and Zhang et al. (2021), which iteratively enhance the labeled dataset through self-training. Although pseudo-labeling methods can effectively improve model performance, they are often affected by the quality of pseudo-labels and can easily lead to model overfitting. The study by Mishra et al. (2024) explores the issue of overconfidence caused by pseudo-labels in semi-supervised learning, analyzes the source of calibration errors, and aims to improve the reliability of the model by evaluating and potentially improving methods.
Consistency Regularization methods improve the generalization ability of models by forcing them to maintain consistency on unlabeled data. The Virtual Adversarial Training (VAT) method proposed by Miyato et al. (2018) enhances the robustness of the model by perturbing the input data and maintaining the stability of the prediction results. The Mean Teacher method proposed by Tarvainen & Valpola (2017) also further improves model performance by using the consistency between the predictions of the teacher model and the output of the student model. Although these methods have achieved good results in many tasks, they often rely on complex training strategies and architectural designs, and their generalization ability across different datasets still needs to be improved.
Semi-Supervised Gaussian Mixture Model (SGMM) is the basic probabilistic framework of our method, combining GMM with unlabeled data, and proposes a learning method that simultaneously considers labeled data and unlabeled data by maximizing the joint log-likelihood function. Miller & Uyar (1996) proposed a classifier structure and learning algorithm that can effectively use unlabeled data to improve performance. There are also many impressive works such as Zong et al. (2018) using deep autoencoders and Gaussian mixture models for anomaly detection.
Most related to ours is the model of Zhao et al. (2023), which also uses ViT for image embedding and then applies GMM for classification. This model is very similar to ViTSGMM, but it addresses the Generalized Category Discovery (GCD) problem from Vaze et al. (2022), aiming to discover new categories, whereas our work goes further to demonstrate that combining deep learning with SGMM can outperform state-of-the-art methods, especially in scenarios with extremely limited labeled data, yielding significant advantages.
2.2 Deep neural networks (DNN) and Visual Transformers (ViT)
Deep Neural Networks (DNNs) have made significant progress in various supervised learning tasks in recent years. The breakthrough achieved by AlexNet, proposed by Krizhevsky et al. (2012), in image classification tasks laid the foundation for the development of deep learning. To improve the generalization ability of DNNs, many methods have been proposed to address overfitting issues, such as Dropout by Srivastava et al. (2014) and BatchNormalization by Ioffe & Szegedy (2015). The Residual Network (ResNet) proposed by He et al. (2016) effectively solved the gradient vanishing problem in deep network training by introducing residual connections, making it possible to train deeper networks. Subsequently, DLA by Yu et al. (2018) introduced a dynamic hierarchical attention mechanism on top of ResNet, further optimizing layer-wise interaction and feature fusion.
Visual Transformers (ViT) are image classification models based on the Transformer architecture that capture global information in images through self-attention mechanisms. Dosovitskiy et al. (2021) proposed dividing images into fixed-size patches and applying self-attention mechanisms. In this way, ViT demonstrates strong performance in image classification tasks, especially on large-scale datasets, exhibiting higher generalization ability compared to traditional CNN architectures. The success of ViT has inspired many researchers to try applying it to self-supervised learning. The DINO method proposed by Caron et al. (2021) learns useful feature representations through self-supervised pre-training of ViT models. These representations can be fine-tuned on a small amount of labeled data to achieve high classification accuracy.
2.3 Self-Supervised Learning and Data Leakage Issues
Self-Supervised Learning is an unsupervised learning approach that automatically learns feature representations through pretext tasks (e.g., image jigsaw puzzles, predicting future frames). The CLIP model proposed by Radford et al. (2021) achieves strong image understanding capabilities through self-supervised learning of image-text contrast, making its performance on unlabeled data very prominent.
In addition, methods like SimCLR from Chen et al. (2020) and MoCo from He et al. (2020) train neural networks through contrastive learning, enabling them to learn discriminative features without labeled data. These methods provide higher quality representations for feature extractors, thereby improving the performance of downstream tasks (such as classification, face recognition, etc.).
In machine learning, the issue of data leakage has always been a potential challenge. Off-topic images, near-duplicate images, and labeling errors in benchmark datasets can lead to inaccurate estimations of model performance. Handling and removing duplicate samples has become an issue that cannot be ignored in semi-supervised learning. Therefore, Gröger et al. (2024) re-examines the task of data cleaning and formalizes it as a ranking problem.
3 Method
We assume that the entire dataset $\mathcal{X}=\{\mathbf{x}_{i}\in\mathbb{R}^{H\times W\times C},i=1,\ldots,N\}$ consists of a labeled dataset $\mathcal{X}_{l}=\{(\mathbf{x}_{1},c_{1}),\ldots,(\mathbf{x}_{N_{l}},c_{N_{l}})\}$ and an unlabeled dataset $\mathcal{X}_{u}=\{\mathbf{x}_{N_{l}+1},\ldots,\mathbf{x}_{N}\}$. Here, the number of labeled data $N_{l}$ is much smaller than the total data amount $N$ ($N_{l}\ll N$).
Our method is as follows:
3.1 Powerful Feature Extractor
For each input image $\mathbf{x}_{i}\in\mathcal{X}$, where $\mathbf{x}_{i}\in\mathbb{R}^{H\times W\times C}$, we divide it into a sequence of fixed-size $P\times P$ patches. Let $N_{p}=\frac{HW}{P^{2}}$ denote the number of patches per image. Each patch $\mathbf{p}^{(i)}_{k}\in\mathbb{R}^{P^{2}C}$ (where $k=1,\ldots,N_{p}$) is flattened into a vector. Then, through a linear transformation, each patch vector $\mathbf{p}^{(i)}_{k}$ is mapped to a $D$-dimensional vector $\mathbf{z}^{(i)}_{k}\in\mathbb{R}^{D}$, expressed as:
$\mathbf{z}^{(i)}_{k}=\mathbf{W}\mathbf{p}^{(i)}_{k}+\mathbf{b}$ (1)
where $\mathbf{W}\in\mathbb{R}^{D\times P^{2}C}$ and $\mathbf{b}\in\mathbb{R}^{D}$ are learnable parameters. A learnable positional embedding $\mathbf{e}_{k}\in\mathbb{R}^{D}$ is added to each patch vector $\mathbf{z}^{(i)}_{k}$ to retain the positional information of the image patches, yielding $\mathbf{z}^{\prime(i)}_{k}=\mathbf{z}^{(i)}_{k}+\mathbf{e}_{k}$. This results in the image patch sequence $\mathbf{Z}^{(i)}=\{\mathbf{z}^{\prime(i)}_{1},\mathbf{z}^{\prime(i)}_{2},\ldots,\mathbf{z}^{\prime(i)}_{N_{p}}\}$.
Subsequently, the image patch sequence $\mathbf{Z}^{(i)}$ is input into a Transformer encoder. The encoder utilizes the self-attention mechanism, which is defined as:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{D}}\right)\mathbf{V}$ (2)
where $\mathbf{Q}=\mathbf{Z}^{(i)}\mathbf{W}_{Q}$, $\mathbf{K}=\mathbf{Z}^{(i)}\mathbf{W}_{K}$, and $\mathbf{V}=\mathbf{Z}^{(i)}\mathbf{W}_{V}$. Through multiple layers of Transformer encoding and Principal Component Analysis (PCA) dimensionality reduction, we finally obtain high-quality image feature vectors $\mathbf{f}_{i}\in\mathbb{R}^{d}$ where $d<D$.
3.2 Semi-supervised Mixture of Experts Classifier
Each feature vector $\mathbf{f}_{i}\in\mathbb{R}^{d}$ is obtained from the feature extractor. Each labeled sample $(\mathbf{f}_{i},c_{i})$ is associated with a class label $c_{i}$ belonging to the label set =$\mathcal{L}=\{1,2,\ldots,K\}$, where $K=\max\{c_{1},\ldots,c_{N_{l}}\}$ represents the total number of classes.
Figure 2: Overview of our ViTSGMM network. The network consists of a feature extractor and a semi-supervised mixture of experts classifier. An example of a sub-model is shown in the ellipse at the bottom right, where the circle labeled 1 represents a feature vector with a corresponding label, 2 represents a pseudo-labeled feature vector, and 3 represents an unlabeled feature vector.
The classifier aims to maximize the following joint log-likelihood function:
$\mathcal{L}(\Theta)=\underbrace{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\log\sum_{l=1}^{L}\pi_{l}p(\mathbf{f}_{i}|\theta_{l})}_{\text{unlabeled data likelihood}}+\underbrace{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\log\sum_{l=1}^{L}\pi_{l}p(\mathbf{f}_{i}|\theta_{l})P(c_{i}|l)}_{\text{labeled data likelihood}}$ (3)
where $\Theta=\{\pi_{1},\ldots,\pi_{L},\theta_{1},\ldots,\theta_{L}\}$ is the set of model parameters with $\sum_{l=1}^{L}\pi_{l}=1$, $\pi_{l}$ represents the mixing coefficient of the $l$-th Gaussian component, $p(\mathbf{f}_{i}|\theta_{l})$ is the probability density function of the $l$-th Gaussian component with parameters $\theta_{l}=\{\bm{\mu}_{l},\Sigma_{l}\}$, and $P(c_{i}|l)$ represents the conditional probability that a feature belongs to class $c_{i}$ given Gaussian component $l$.
The EM algorithm updates parameters as follows:
$\gamma_{il}^{(t)}=\frac{\pi_{l}^{(t)}p(\mathbf{f}_{i}|\theta_{l}^{(t)})}{\sum_{j=1}^{L}\pi_{j}^{(t)}p(\mathbf{f}_{i}|\theta_{j}^{(t)})}\quad\forall\mathbf{f}_{i}\in\mathcal{F}_{u}$ (4)
$\gamma_{il|c_{i}}^{(t)}=\frac{\pi_{l}^{(t)}p(\mathbf{f}_{i}|\theta_{l}^{(t)})P(c_{i}|l)^{(t)}}{\sum_{j=1}^{L}\pi_{j}^{(t)}p(\mathbf{f}_{i}|\theta_{j}^{(t)})P(c_{i}|j)^{(t)}}\quad\forall\mathbf{f}_{i}\in\mathcal{F}_{l}$ (5)
$\bm{\mu}_{l}^{(t+1)}=\frac{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\mathbf{f}_{i}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\mathbf{f}_{i}\gamma_{il}^{(t)}}{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\gamma_{il}^{(t)}}$ (6)
$\Sigma_{l}^{(t+1)}=\frac{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}M_{il}^{(t)}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}M_{il}^{(t)}\gamma_{il}^{(t)}}{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\gamma_{il}^{(t)}}$ (7)
where $M_{il}^{(t)}=(\mathbf{f}_{i}-\bm{\mu}_{l}^{(t)})(\mathbf{f}_{i}-\bm{\mu}_{l}^{(t)})^{\top}$.
$\pi_{l}^{(t+1)}=\frac{1}{N}\left(\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\gamma_{il}^{(t)}\right)$ (8)
$P(k|l)^{(t+1)}=\frac{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l},c_{i}=k}\gamma_{il|c_{i}}^{(t)}}{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\gamma_{il|c_{i}}^{(t)}}$ (9)
3.3 Pseudo-labeling Mechanism
To effectively leverage information from the unlabeled dataset $\mathcal{X}_{u}$, we introduce a pseudo-labeling mechanism based on the component-class association probabilities. Unlike traditional iterative methods, our pseudo-label generation is performed in a single calculation after the initial SGMM training reaches convergence, reducing computational cost and mitigating error accumulation risks.
For unlabeled features $\mathbf{f}_{i}\in\mathcal{F}_{u}$, we compute class probabilities through Gaussian component responsibilities:
$p_{i}^{(k)}=\sum_{l=1}^{L}P(k|l)^{(t)}\gamma_{il}^{(t)},\quad\xi_{i}=\max_{1\leq k\leq K}p_{i}^{(k)}$ (10)
where $P(k|l)$ represents the class probability conditioned on component $l$, and $\gamma_{il}^{(t)}$ is the responsibility from Equation 4.
We construct class-specific candidate sets with confidence thresholding sorted by $\xi_{i}$ descending:
$\mathcal{C}_{k}=\left\{\mathbf{f}_{i}\mid\arg\max_{k}p_{i}^{(k)}=k,\ \xi_{i}>\tau,\quad\tau\in(0,1)\right\}$ (11)
To prevent class imbalance, we adopt proportional sampling:
$n^{\text{final}}=\min_{1\leq k\leq K}\lfloor\alpha|\mathcal{C}_{k}|\rfloor\quad(\alpha\in(0,1))$ (12)
The pseudo-labeled set is then constructed as:
$\mathcal{D}_{p}=\{(\mathbf{f}_{i},k)\mid k=1,\ldots,K,\ \mathbf{f}_{i}\in\mathcal{C}_{k}[1:n^{\text{final}}]\}$ (13)
Finally, combine the pseudo-labeled data $\mathcal{D}_{p}$ with the labeled features $\mathcal{F}_{l}$ and use the EM algorithm for iterative. We use Kmeans++ to initialize the parameters first, during the training process, log-likelihood value calculated using Equation 3 increases steadily. We show the procedure of our method in Algorithm 1.
Algorithm 1 ViTSGMM Network
0:  Labeled data $\mathcal{X}_{l}$, Unlabeled data $\mathcal{X}_{u}$, Pre-trained ViT, SGMM components $L$, Confidence threshold $\tau$, Sampling ratio $\alpha$
0:  Trained SGMM parameters $\Theta$
1:  Extract features $\mathcal{F}_{l}\leftarrow\{\text{ViT}(\mathbf{x}_{i})|(\mathbf{x}_{i},c_{i})\in\mathcal{X}_{l}\}$ and $\mathcal{F}_{u}\leftarrow\{\text{ViT}(\mathbf{x}_{i})|\mathbf{x}_{i}\in\mathcal{X}_{u}\}$
2:  PCA $\mathcal{F}_{l}$ and $\mathcal{F}_{u}$ into $\mathbb{R}^{d}$
3:  $L$ centers cluster $\mathcal{F}_{l}\cup\mathcal{F}_{u}$ with K-means++
4:  Initialize $\{\pi_{l},\bm{\mu}_{l},\Sigma_{l}\}$ from cluster results
5:  for $t=1$ to $T_{1}$ do
6:     EM algorithm: Initial iterative via (4-9)
7:  end for
8:  Construct $\mathcal{D}_{p}$ using (11-13) with $\tau$ and $\alpha$
9:  Augment labeled set: $\mathcal{F}_{l}\leftarrow\mathcal{F}_{l}\cup\mathcal{D}_{p}$
10:  for $t=1$ to $T_{2}$ do
11:     EM algorithm: Final iterative via (4-9)
12:  end for
13:  return  $\Theta$
4 Experiments
4.1 Standard SGMM Performance Analysis
The performance of the SGMM is influenced by several factors. This section will delve into the impact of the number of labels, feature dimensions, and the number of Gaussian components on the model’s behavior.
SGMM’s core of its labeling efficiency lies in using unlabeled data to assist the model in learning the data distribution. Theoretically, as the number of labeled samples increases, the model can more accurately estimate the conditional probability of categories $P(k|l)$, thereby improving discrimination ability. When the labeled samples reach a certain scale, the marginal benefit of the additional information provided by unlabeled data diminishes, and the performance improvement curve tends to flatten. Therefore, SGMM’s advantage is that it can quickly improve performance with very few labels and demonstrate high efficiency in applications with limited labeling resources.
Figure 3: Relationship between the number of labels and performance. The trend of accuracy with the number of labels for SGMM on the MNIST dataset shows a steady improvement in accuracy with an increase in labels.
When dealing with high-dimensional data, the feature dimension is crucial to the performance and computational efficiency of SGMM. High-dimensional feature spaces can lead to the “curse of dimensionality”, increasing the complexity of model training and the risk of overfitting, while also consuming excessive computational resources. Reducing feature dimensions through dimensionality reduction techniques such as PCA can retain discriminative feature information while removing redundant dimensions, alleviating these problems to some extent. However, excessive dimensionality reduction can lead to the loss of key feature information. Therefore, the feature dimension should be reduced as much as possible while ensuring model performance.
Figure 4: Relationship between PCA dimensions/Gaussian components and performance. The left figure shows the trend of SGMM accuracy with PCA dimensions on the MNIST dataset, while the right figure shows the trend of accuracy with the number of Gaussian submodels.
Gaussian Mixture Models use multiple Gaussian components to fit complex data distributions, and the number of components directly affects the model’s complexity and fitting ability. When the number of components is small, the model may not be able to fully capture the fine-grained structure of the data distribution, leading to underfitting. When the number of components is too large, the model complexity increases, which can easily lead to overfitting of the training data, reducing generalization ability and increasing computational costs. Therefore, the optimal number of Gaussian components should match the intrinsic class structure and complexity of the dataset.
In summary, labeling efficiency reflects the core advantage of semi-supervised learning. Feature dimension efficiency emphasizes the role of dimensionality reduction in improving model efficiency and generalization ability. The number of Gaussian components reflects the trade-off between model complexity and data fitting ability. Understanding the impact of these factors on SGMM performance helps to adjust model parameters according to specific tasks in practical applications.
4.2 More Powerful Feature Extractors
To further analyze the performance of the SGMM, we experimented on the CIFAR-10 dataset using different feature extractors: PCA, Deep Neural Networks (ResNet101, DLA169), and ViT-Base/Large backbone DINO pre-trained models. Figure 5 shows that with DNNs and ViTs, increasing the number of Gaussian components leads to a decrease in classification accuracy, which contrasts with the performance trend observed when using PCA.
Figure 5: Performance trends. The left figure shows the accuracy trend of SGMM on the CIFAR-10 dataset processed with PCA, while the right figure shows the accuracy trend with the number of Gaussian submodels when using DNNs and ViTs.
Next, we performed t-SNE visualization on the features extracted by the three methods. In Figure 6, we present the t-SNE visualization results of feature representations obtained using different feature extraction methods on the CIFAR-10 dataset. It can be clearly seen from the figure that the feature points extracted by PCA exhibit a highly mixed state in the t-SNE space, and it is difficult to distinguish between samples from different categories. This indicates that the feature representation after PCA dimensionality reduction lacks sufficient discriminative power. In contrast, the feature points extracted using deep neural networks exhibit a certain clustering trend in the t-SNE space, but there is still some overlap between the boundaries of different categories. The feature points extracted using Visual Transformer (ViT) form a relatively clear cluster structure in the t-SNE space, which indicates that the feature representation extracted by ViT has stronger separability and can better capture the intrinsic structure of the data, providing a more favorable feature basis for subsequent classification tasks.
PCA has limited feature discrimination ability, so it needs more Gaussian sub-models to model complex intra-class variations. At this time, increasing the number of Gaussian components helps to approximate the true distribution; while the features extracted by DNN/ViT are close to a single Gaussian for each category, increasing the number of components will cause the model complexity to exceed the actual needs. These analyses are consistent with the experimental results in Figure 5.
Figure 6: t-SNE Visualization. t-SNE visualization results of different feature extraction methods on the CIFAR-10 dataset.
We analyzed high-confidence samples from two Gaussian components of the same bird category. Observing the original images in Figure 7, we found that these two Gaussian components exhibit different feature preferences: one component tends to give high confidence to samples of ostriches (long-necked birds), while the other component prefers plump, neckless birds.
Figure 7: Comparison raw images. Visualization of high-confidence samples from two Gaussian components of the bird category on the CIFAR-10 dataset.
Through experimental results and visualization analysis, we found that DNN and ViT can better capture the distribution characteristics of data compared to the traditional PCA dimensionality reduction method, thereby improving the classification performance of SGMM.
4.3 Experiment Setup
Datasets.
We report results on the Cifar-10/100 and STL-10 datasets. CIFAR-10 contains 10 classes with 6,000 32x32 color images per class, totaling 60,000 images, of which 50,000 are for training and 10,000 for testing. CIFAR-100 is similar to CIFAR-10 but contains 100 classes with 600 images per class, of which 500 are for training and 100 for testing. STL-10 is an image classification dataset designed for self-supervised and semi-supervised learning, with 10 classes in the training set, containing 5,000 labeled samples and 100,000 unlabeled images, where images have a resolution of 96x96 pixels. We performed special deduplication processing on the STL-10 dataset, remains 90,455 training images and 8,000 images for testing.
Implementation details.
We use the DINO pre-trained model based on the ViT-large backbone to extract feature representations. On the Cifar-10/100 datasets, the values of the Gaussian components are set to 10/100 respectively, and the PCA dimension is set to 60. In the STL-10 dataset, since its training set contains more than 10 categories, we set the number of Gaussian components to 15 to obtain better robustness, the PCA dimension is set to 45, and the convergence threshold (Tol) is set to 1. The PCA dimension is set to use accuracy as an evaluation metric to evaluate the classification performance of different methods. All experiments are performed on a Tesla T4 GPU with 15GB memory.
4.4 Comparison with State-of-the-art Methods
Results on CIFAR-10.
We show the benchmark results on CIFAR-10 in Table 1. Our method achieves competitive performance compared with state-of-the-art methods, including FixMatch, UDA, and ReMixMatch. It can be clearly seen that our method consistently outperforms other methods under all settings.
Table 1: Benchmark results on CIFAR-10.
Algorithms Error Rate (40 labels) Error Rate (250 labels) Error Rate (4000 labels)
Dash (2021) 9.29 $\pm$ 3.28 5.16 $\pm$ 0.23 4.36 $\pm$ 0.11
MPL (2021) 6.62 $\pm$ 0.91 5.76 $\pm$ 0.24 4.36 $\pm$ 0.11
FlexMatch (2021) 4.97 $\pm$ 0.06 4.98 $\pm$ 0.09 4.19 $\pm$ 0.01
CoMatch (2021) 6.51 $\pm$ 1.18 5.35 $\pm$ 0.14 4.27 $\pm$ 0.12
SimMatch (2022) 5.38 $\pm$ 0.01 5.36 $\pm$ 0.08 4.41 $\pm$ 0.07
AdaMatch (2022) 5.09 $\pm$ 0.21 5.13 $\pm$ 0.05 4.36 $\pm$ 0.05
FreeMatch (2023) 4.90 $\pm$ 0.12 4.88 $\pm$ 0.09 4.16 $\pm$ 0.06
SoftMatch (2023) 5.11 $\pm$ 0.14 4.96 $\pm$ 0.09 4.27 $\pm$ 0.05
SequenceMatch (2024a) 4.80 $\pm$ 0.01 4.75 $\pm$ 0.05 4.15 $\pm$ 0.01
EPASS (2024b) 5.55 $\pm$ 0.21 5.31 $\pm$ 0.13 4.23 $\pm$ 0.05
ViTSGMM(Ours) 3.51 $\pm$ 0.12 3.47 $\pm$ 0.17 3.45 $\pm$ 0.16
Results on STL-10.
In Table 2, we present the performance on STL-10. Our method achieves state-of-the-art performance on STL-10, outperforming other methods by a large margin. Our method achieves the best performance in all three settings, demonstrating the effectiveness of our proposed method. Compared to SequenceMatch, our ViTSGMM achieves +10.88%, +8.35%, and +1.41% on 40-label, 250-label, and 1000-label settings, respectively.
Table 2: Benchmark results on STL-10.
Algorithms Error Rate (40 labels) Error Rate (250 labels) Error Rate (1000 labels)
Dash (2021) 42.00 $\pm$ 4.94 10.50 $\pm$ 1.37 6.30 $\pm$ 0.49
MPL (2021) 35.97 $\pm$ 4.14 9.90 $\pm$ 0.96 6.66 $\pm$ 0.00
FlexMatch (2021) 29.12 $\pm$ 5.04 9.85 $\pm$ 1.35 6.08 $\pm$ 0.34
CoMatch (2021) 13.74 $\pm$ 4.20 7.63 $\pm$ 0.94 5.71 $\pm$ 0.08
SimMatch (2022) 16.98 $\pm$ 4.24 8.27 $\pm$ 0.40 5.74 $\pm$ 0.31
AdaMatch (2022) 19.95 $\pm$ 5.17 8.59 $\pm$ 0.43 6.01 $\pm$ 0.02
FreeMatch (2023) 28.50 $\pm$ 5.41 9.29 $\pm$ 1.24 5.81 $\pm$ 0.32
SoftMatch (2023) 22.23 $\pm$ 3.82 9.18 $\pm$ 0.63 5.79 $\pm$ 0.15
SequenceMatch (2024a) 15.45 $\pm$ 1.40 12.78 $\pm$ 0.76 5.56 $\pm$ 0.35
EPASS (2024b) 9.15 $\pm$ 3.25 6.27 $\pm$ 0.03 5.40 $\pm$ 0.12
ViTSGMM(Ours) 4.57 $\pm$ 0.24 4.43 $\pm$ 0.08 4.15 $\pm$ 0.07
Results on CIFAR-100.
In Table 3, we compare the performance of ViTSGMM with state-of-the-art methods on CIFAR-100 dataset. Our method outperforms other methods in the 40 labels setting, and achieves comparable performance in the 2500 labels and 10,000 labels settings.
Table 3: Benchmark results on CIFAR-100.
Algorithms Error Rate (400 labels) Error Rate (2500 labels) Error Rate (10000 labels)
Dash (2021) 44.82 $\pm$ 0.96 27.15 $\pm$ 0.22 21.88 $\pm$ 0.07
MPL (2021) 46.26 $\pm$ 1.84 27.71 $\pm$ 0.19 21.74 $\pm$ 0.09
FlexMatch (2021) 39.94 $\pm$ 1.62 26.49 $\pm$ 0.20 21.90 $\pm$ 0.15
CoMatch (2021) 53.41 $\pm$ 2.36 29.78 $\pm$ 0.11 22.11 $\pm$ 0.22
SimMatch (2022) 39.32 $\pm$ 0.72 26.21 $\pm$ 0.37 21.50 $\pm$ 0.11
AdaMatch (2022) 38.08 $\pm$ 1.35 26.66 $\pm$ 0.33 21.99 $\pm$ 0.15
FreeMatch (2023) 39.52 $\pm$ 0.01 26.22 $\pm$ 0.08 21.81 $\pm$ 0.17
SoftMatch (2023) 37.60 $\pm$ 0.24 26.39 $\pm$ 0.38 21.86 $\pm$ 0.16
SequenceMatch (2024a) 37.86 $\pm$ 1.07 25.99 $\pm$ 0.22 20.10 $\pm$ 0.04
EPASS (2024b) 38.88 $\pm$ 0.24 25.68 $\pm$ 0.33 21.32 $\pm$ 0.14
ViTSGMM(Ours) 26.59 $\pm$ 1.02 22.19 $\pm$ 0.81 21.21 $\pm$ 0.26
4.5 Ablation Study
Comparing Semi-supervised Learning Methods.
In order to test the effectiveness of the SGMM component, we compared it with the original classification head of ViT.
We trained on three datasets, error rates shown in Table 4, under scenarios with extremely few labels, the performance of SGMM is significantly superior to that of the Softmax layer.
Effect of Pseudo-labeling.
We conducted an ablation study to evaluate the impact of pseudo-labeling on the performance of ViTSGMM. We compared the performance of ViTSGMM with and without pseudo-labeling on three datasets, error rates shown in Table 4.
Table 4: Ablation study with three datasets.
Dataset CIFAR-10 CIFAR-100 STL-10
# Label 40 250 4000 400 2500 10000 40 250 1000
w Softmax 51.05 $\pm$ 16.69 5.40 $\pm$ 1.70 3.75 $\pm$ 0.21 57.32 $\pm$ 10.91 43.01 $\pm$ 5.89 33.70 $\pm$ 2.71 16.29 $\pm$ 2.85 4.88 $\pm$ 1.50 4.38 $\pm$ 0.40
w/o P-L 3.73 $\pm$ 0.21 3.70 $\pm$ 0.23 3.63 $\pm$ 0.19 30.56 $\pm$ 0.80 23.21 $\pm$ 0.66 21.56 $\pm$ 0.43 4.91 $\pm$ 0.12 4.81 $\pm$ 0.12 4.58 $\pm$ 0.10
ViTSGMM 3.51 $\pm$ 0.12 3.47 $\pm$ 0.17 3.45 $\pm$ 0.16 26.59 $\pm$ 1.02 22.19 $\pm$ 0.81 21.21 $\pm$ 0.26 4.44 $\pm$ 0.04 4.43 $\pm$ 0.08 4.15 $\pm$ 0.07
5 Conclusions
We introduce high-performance feature extractors into traditional generative models. Unlike traditional semi-supervised learning methods, this approach does not introduce domain-specific constraints or assumptions. The cross-modal unified modeling strategy performs surprisingly well when trained with a very small amount of labeled data and a large amount of unlabeled data. Extensive experiments show that ViTSGMM achieves new performance breakthroughs in benchmarks such as STL-10 and Cifar-10, maintaining over 96% classification accuracy even when using only 4 labeled samples per class.
Future development of this technology can proceed along several dimensions. First, constructing dynamic component adjustment mechanism based on a Bayesian framework to achieve parameter adaptivity. Secondly, enhancing the model’s discriminative performance when facing complex tasks such as fine-grained classification. Most interesting, exploring its application in other fields, such as electroencephalogram (EEG) signal analysis and medical image processing, thereby broadening its application scope and verifying its transferability. We believe these improvements will further unlock the technological benefits brought about by the fusion of probabilistic graphical models and deep learning.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62076078 and in part by the Chinese Association for Artificial Intelligence (CAAI)-Huawei MindSpore Open Fund under Grant CAAIXSJLJJ-2020-033A.
Appendix
Appendix A STL-10 data clean
Duplicate samples in the STL-10 dataset can potentially bias model evaluation. To address this, we deduplicated the dataset. Specifically, we used a script to compute image hashes for STL-10 dataset, then removed 7,545 samples from the trainset that were found to be duplicates of samples in the testset. The pseudo-code for this deduplication process is provided in Algorithm A.1.
Algorithm A.1 Deduplication of Dataset
0:  Main dataset, test dataset, batch size
0:  Deduplicated dataset, CSV file with duplicate information
  Build test set hash dictionary: Compute hash for each test image
2:  Initialize valid_indices and duplicates as empty lists
  for each batch in main dataset do
4:     for each image in batch do
        Compute hash of the image
6:        if hash exists in test set hash dictionary then
           Record duplicate sample details
8:        else
           Add current index to valid_indices
10:        end if
     end for
12:  end for
  if duplicates is not empty then
14:     Save duplicate details to CSV file
  end if
Following processing, the training set consists of 5,000 labeled samples and 90,455 unsupervised samples. This effectively addresses the problem of data leakage, thereby guaranteeing the credibility of the experimental findings.
Figure B.1: Heatmap comparison. From left to right: raw image, DLA169, DINO. Note the small penguin on the right in the third row.
Appendix B Comparing Heatmaps of DNN and ViT
We compare the heatmaps of DNN and ViT on several randomly selected images from the internet, each containing two animals simultaneously. The heatmaps are generated using the Grad-CAM method, which visualizes the importance of each pixel in the image for the classification result. As shown in Figure B.1, the heatmaps of DNN and ViT differ significantly. While DNN focuses primarily on one animal, ViT attends to both animals in the image, demonstrating its ability to capture global context more effectively. This comparison highlights the differences in how these models prioritize regions of the image during classification.
Appendix C PCA Explained Variance
In our experiments, the hyperparameter of PCA dimensions for different datasets is determined by the PCA explained variance, which represents the percentage of variance in the data that can be explained by the feature vectors after PCA dimensionality reduction. We typically choose dimensions that can explain over 60% of the variance.
Figure C.1: PCA explained variance. The percentage of variance explained by each PCA dimension on the CIFAR-10 dataset.
Appendix D ViT Classification Head for Semi-Supervised Learning
The following is the mathematical implementation of the original ViT classification head for semi-supervised learning, which has been used in our ablation study in section 4.5.
In section 3.1, we define the feature vector $\mathbf{f}_{i}\in\mathbb{R}^{d}$. The standard ViT classification head includes a linear transformation $\mathbf{s}_{i}=\mathbf{W}_{h}\mathbf{f}_{i}+\mathbf{b}_{h}$, where $\mathbf{W}_{h}\in\mathbb{R}^{K\times d}$ and $\mathbf{b}_{h}\in\mathbb{R}^{K}$ are learnable parameters, and $K$ is the number of categories; followed by softmax:
$P(c=k|\mathbf{f}_{i})=\frac{\exp(s_{i,k})}{\sum_{j=1}^{K}\exp(s_{i,j})},\quad\forall k\in\{1,\ldots,K\}$
For labeled samples $\mathcal{F}_{l}=\{(\mathbf{f}_{i},c_{i})\}$, we optimize the standard cross-entropy loss:
$\mathcal{L}_{\text{sup}}=-\frac{1}{|\mathcal{F}_{l}|}\sum_{(\mathbf{f}_{i},c_{i})\in\mathcal{F}_{l}}\log P(c=c_{i}|\mathbf{f}_{i})$
Unlabeled data $\mathcal{F}_{u}=\{\mathbf{f}_{j}\}$ adopts a confidence-based pseudo-labeling strategy, where the pseudo-label is $\hat{c}_{j}=\arg\max_{k}P(c=k|\mathbf{f}_{j})$, and the confidence score is $\xi_{j}=\max_{k}P(c=k|\mathbf{f}_{j})$. The unsupervised loss is calculated as follows:
$\mathcal{L}_{\text{unsup}}=-\frac{1}{|\mathcal{F}_{u}|}\sum_{\mathbf{f}_{j}\in\mathcal{F}_{u}}\mathbb{I}(\xi_{j}>\tau)\log P(c=\hat{c}_{j}|\mathbf{f}_{j})$
The total loss combines the supervised and unsupervised objectives, with the weight $\lambda(t)$ changing over time, where $T_{\text{ramp}}$ corresponds to the training period:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{sup}}+\underbrace{\lambda_{\text{max}}\cdot\min\left(1,\frac{t}{T_{\text{ramp}}}\right)}_{\lambda(t)}\mathcal{L}_{\text{unsup}}$