### v2

# SemiOccam: A Robust Semi-Supervised Image Recognition Network Using Sparse Labels

Abstract
We present SemiOccam, an image recognition network that leverages semi-supervised learning in a highly efficient manner. Existing works often rely on complex training techniques and architectures, requiring hundreds of GPU hours for training, while their generalization ability when dealing with extremely limited labeled data remains to be improved. To address these limitations, we construct a hierarchical mixture density classification decision mechanism by optimizing mutual information between feature representations and target classes, compressing redundant information while retaining crucial discriminative components. Experimental results demonstrate that our method achieves state-of-the-art performance on various datasets when using negligible labeled samples, and its simple architecture keeps training time to minute-level. Notably, this paper reveals a long-overlooked data leakage issue in the STL-10 dataset for semi-supervised learning tasks and removes duplicates to ensure the reliability of experimental results. We also release the deduplicated CleanSTL-10 dataset to facilitate fair and reliable research in future semi-supervised learning.  [^1] Code available at https://github.com/Shu1L0n9/SemiOccam.  [^2] CleanSTL-10 available at https://huggingface.co/datasets/Shu1L0n9/CleanSTL-10.
Keywords: Semi-Supervised Learning, Image Recognition, Vision Transformer, Gaussian Mixture Model
1 Introduction
Deep learning has achieved remarkable success in image classification tasks. However, its performance is highly dependent on large-scale annotated datasets. In real-world applications, labeled data is often scarce and expensive to acquire. Therefore, semi-supervised learning (SSL) has received widespread attention in recent years as an effective solution. The goal of semi-supervised learning is to train models using a small amount of labeled data and large amount of unlabeled data, thereby improving the model’s generalization ability. Nevertheless, most existing SSL methods, such as consistency regularization and pseudo-labeling, rely on complex training strategies and network designs, leading to training costs of hundreds of GPU hours even on high-performance computing resources like NVIDIA V100, and their generalization ability when labeled data are extremely limited still needs improvement. Moreover, although some recent works have explored Vision Transformers (ViT) in SSL, they have yet to fully exploit its advantages.
Gaussian Mixture Models (GMMs) are powerful probabilistic models that assume data is generated from a mixture of multiple Gaussian distributions. The inspiration for this research comes from Miller & Uyar’s systematic study of the “Mixture of Experts (MoE)” structure based on the joint probability model of features and labels, as well as the principle of Occam’s razor.
According to Occam’s razor, among competing hypotheses that explain the data well, the simplest one is preferred. This motivates us to design a model that balances complexity and interpretability, ensuring efficient and effective feature representation without unnecessary overfitting. In our research, we found that the performance of GMMs is highly related to the quality of feature vectors. If the features lack representativeness or have insufficient discriminative power, the classification performance will be severely affected.
To address this, we designed a feature encoding framework based on a self-attention mechanism to fully exploit the global contextual information in images and generate more discriminative feature representations. Simultaneously, we incorporated semi-supervised learning strategies, utilizing a pseudo-labeling mechanism to efficiently leverage unlabeled data.
This approach achieves superior classification performance within the framework of mixture density estimation, significantly overcomes the limitations of traditional methods in terms of feature representation capability, training cost and model generalization performance. Taking the Dino with ViT-large backbone as our feature extraction model, we validated the proposed semi-supervised Gaussian mixture model (SGMM) method on various datasets, results show that our method achieves state-of-the-art performance within minutes on a single Tesla T4 GPU and exhibits good stability and robustness, with strong potential for edge deployment.
Figure 1: Performance Comparison. Comparison of our method with recent semi-supervised learning models on datasets with different numbers of labels, achieving state-of-the-art performance with low training cost.
It is worth mentioning that we have revealed a serious problem that has been overlooked by existing semi-supervised learning research: there are 7,500+ duplicate samples between the training samples and the testing samples in STL-10, which is undoubtedly a serious data leakage problem when directly used for semi-supervised learning tasks. To address this, we release the deduplicated CleanSTL-10 dataset for fair and reliable evaluation.
Our main contributions can be summarized as follows:
  • We introduce the powerful ViT as a feature extractor into the classical yet often underappreciated MoE classifier for image recognition tasks.
  • We design a simple but performant semi-supervised learning strategy that effectively leverages negligible labels, achieves state-of-the-art performance.
  • We reveal the longstanding, yet overlooked issue of training set and test set contamination in the STL-10 dataset, which has not been addressed in other works.
  • We release the deduplicated CleanSTL-10 dataset to promote fair and reliable future research.
The remainder of this paper is organized as follows: Section 2 reviews related works; Section 3 introduces our SemiOccam network; Section 4 demonstrates and analyzes experimental results, including both classification performance and training cost; Section 5 discusses the significance and limitations of the research findings and outlines future research directions.
2 Related Works
2.1 Semi-Supervised Learning and Gaussian Mixture Models
Semi-Supervised Learning has been widely applied in tasks such as image classification and speech recognition in recent years, especially in scenarios where labeled data is scarce. Early SSL methods mainly relied on generating pseudo-labels and consistency regularization.
Pseudo-Labeling is one of the classic methods in SSL, with the core idea of using an existing model to predict unlabeled data Yarowsky (1995) and adding its predicted labels to the training set as pseudo-labels. Representative works include Lee et al. (2013) and Zhang et al. (2021), which iteratively enhance the labeled dataset through self-training. Although pseudo-labeling methods can effectively improve model performance, they are often affected by the quality of pseudo-labels and can easily lead to model overfitting. The study by Mishra et al. (2024) explores the issue of overconfidence caused by pseudo-labels in semi-supervised learning, analyzes the source of calibration errors, and aims to improve the reliability of the model by evaluating and potentially improving methods.
Consistency Regularization methods improve the generalization ability of models by forcing them to maintain consistency on unlabeled data. The Virtual Adversarial Training (VAT) method proposed by Miyato et al. (2018) enhances the robustness of the model by perturbing the input data and maintaining the stability of the prediction results. The Mean Teacher method proposed by Tarvainen & Valpola (2017) also further improves model performance by using the consistency between the predictions of the teacher model and the output of the student model. Although these methods have achieved good results in many tasks, they often rely on complex training strategies and architectural designs, and their generalization ability across different datasets still needs to be improved.
Semi-Supervised Gaussian Mixture Model is the basic probabilistic framework of our method. It combines GMM with unlabeled data and proposes a learning method that simultaneously considers both labeled and unlabeled data by maximizing the joint log-likelihood function. Our method is inspired by the classifier structure and learning algorithm proposed by Miller & Uyar, which effectively leverages unlabeled data to improve performance. In addition, many impressive works, such as Zong et al. (2018), have extended GMMs by incorporating deep autoencoders for tasks like anomaly detection.
Most related to ours is the model of Zhao et al. (2023), which also uses ViT for image embedding and then applies GMM for classification. This model is very similar to SemiOccam, but it addresses the Generalized Category Discovery (GCD) problem from Vaze et al. (2022), aiming to discover new categories, whereas our work goes further to demonstrate that combining deep learning with SGMM can outperform state-of-the-art methods, especially in scenarios with extremely limited labeled data, yielding significant advantages.
2.2 Deep neural networks and Visual Transformers
Deep Neural Networks have made significant progress in various supervised learning tasks in recent years. The breakthrough achieved by AlexNet, proposed by Krizhevsky et al. (2012), in image classification tasks laid the foundation for the development of deep learning. To improve the generalization ability of Deep Neural Networks (DNNs), many methods have been proposed to address overfitting issues, such as Dropout by Srivastava et al. (2014) and BatchNormalization by Ioffe & Szegedy (2015). The Residual Network (ResNet) proposed by He et al. (2016) effectively solved the gradient vanishing problem in deep network training by introducing residual connections, making it possible to train deeper networks. Subsequently, DLA by Yu et al. (2018) introduced a dynamic hierarchical attention mechanism on top of ResNet, further optimizing layer-wise interaction and feature fusion.
Visual Transformers are image classification models based on the Transformer architecture that capture global information in images through self-attention mechanisms. Dosovitskiy et al. (2021) proposed dividing images into fixed-size patches and applying self-attention mechanisms. In this way, ViT demonstrates strong performance in image classification tasks, especially on large-scale datasets, exhibiting higher generalization ability compared to traditional CNN architectures. The success of ViT has inspired many researchers to try applying it to self-supervised learning. The DINO method proposed by Caron et al. (2021) learns useful feature representations through self-supervised pre-training of ViT models. These representations can be fine-tuned on a small amount of labeled data to achieve high classification accuracy.
2.3 Self-Supervised Learning and Data Leakage Issues
Self-Supervised Learning is an unsupervised learning approach that automatically learns feature representations through pretext tasks (e.g., image jigsaw puzzles, predicting future frames). The CLIP model proposed by Radford et al. (2021) achieves strong image understanding capabilities through self-supervised learning of image-text contrast, making its performance on unlabeled data very prominent.
In addition, methods like SimCLR from Chen et al. (2020) and MoCo from He et al. (2020) train neural networks through contrastive learning, enabling them to learn discriminative features without labeled data. These methods provide higher quality representations for feature extractors, thereby improving the performance of downstream tasks (such as classification, face recognition, etc.).
Data leakage issue has always been a potential challenge in machine learning. Off-topic images, near-duplicate images, and labeling errors in benchmark datasets can lead to inaccurate estimations of model performance. Handling and removing duplicate samples has become an issue that cannot be ignored in semi-supervised learning. Therefore, Gröger et al. (2024) re-examines the task of data cleaning and formalizes it as a ranking problem.
3 Method
We assume that the entire dataset $\mathcal{X}=\{\mathbf{x}_{i}\in\mathbb{R}^{H\times W\times C},i=1,\ldots,N\}$ consists of a labeled dataset $\mathcal{X}_{l}=\{(\mathbf{x}_{1},c_{1}),\ldots,(\mathbf{x}_{N_{l}},c_{N_{l}})\}$ and an unlabeled dataset $\mathcal{X}_{u}=\{\mathbf{x}_{N_{l}+1},\ldots,\mathbf{x}_{N}\}$. Here, the number of labeled data $N_{l}$ is much smaller than the total data amount $N$ ($N_{l}\ll N$).
Our method is as follows:
3.1 Powerful Feature Extractor
For each input image $\mathbf{x}_{i}\in\mathcal{X}$, where $\mathbf{x}_{i}\in\mathbb{R}^{H\times W\times C}$, we divide it into a sequence of fixed-size $P\times P$ patches. Let $N_{p}=\frac{HW}{P^{2}}$ denote the number of patches per image. Each patch $\mathbf{p}^{(i)}_{k}\in\mathbb{R}^{P^{2}C}$ (where $k=1,\ldots,N_{p}$) is flattened into a vector. Then, through a linear transformation, each patch vector $\mathbf{p}^{(i)}_{k}$ is mapped to a $D$-dimensional vector $\mathbf{z}^{(i)}_{k}\in\mathbb{R}^{D}$, expressed as:
$\mathbf{z}^{(i)}_{k}=\mathbf{W}\mathbf{p}^{(i)}_{k}+\mathbf{b}$ (1)
where $\mathbf{W}\in\mathbb{R}^{D\times P^{2}C}$ and $\mathbf{b}\in\mathbb{R}^{D}$ are learnable parameters. A learnable positional embedding $\mathbf{e}_{k}\in\mathbb{R}^{D}$ is added to each patch vector $\mathbf{z}^{(i)}_{k}$ to retain the positional information of the image patches, yielding $\mathbf{z}^{\prime(i)}_{k}=\mathbf{z}^{(i)}_{k}+\mathbf{e}_{k}$. This results in the image patch sequence $\mathbf{Z}^{(i)}=\{\mathbf{z}^{\prime(i)}_{1},\mathbf{z}^{\prime(i)}_{2},\ldots,\mathbf{z}^{\prime(i)}_{N_{p}}\}$.
Subsequently, the image patch sequence $\mathbf{Z}^{(i)}$ is input into a Transformer encoder. The encoder utilizes the self-attention mechanism, which is defined as:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{D}}\right)\mathbf{V}$ (2)
where $\mathbf{Q}=\mathbf{Z}^{(i)}\mathbf{W}_{Q}$, $\mathbf{K}=\mathbf{Z}^{(i)}\mathbf{W}_{K}$, and $\mathbf{V}=\mathbf{Z}^{(i)}\mathbf{W}_{V}$. Through multiple layers of Transformer encoding and Principal Component Analysis (PCA) dimensionality reduction, we finally obtain high-quality image feature vectors $\mathbf{f}_{i}\in\mathbb{R}^{d}$ where $d<D$.
3.2 Semi-supervised Mixture of Experts Classifier
Each feature vector $\mathbf{f}_{i}\in\mathbb{R}^{d}$ is obtained from the feature extractor. Each labeled sample $(\mathbf{f}_{i},c_{i})$ is associated with a class label $c_{i}$ belonging to the label set =$\mathcal{L}=\{1,2,\ldots,K\}$, where $K=\max\{c_{1},\ldots,c_{N_{l}}\}$ represents the total number of classes.
Figure 2: Overview of our SemiOccam network. The network consists of a feature extractor and a semi-supervised mixture of experts classifier. An example of a sub-model is shown in the ellipse at the bottom right, where the circle labeled 1 represents a feature vector with a corresponding label, 2 represents a pseudo-labeled feature vector, and 3 represents an unlabeled feature vector.
The classifier aims to maximize the following joint log-likelihood function:
$\mathcal{L}(\Theta)=\underbrace{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\log\sum_{l=1}^{L}\pi_{l}p(\mathbf{f}_{i}|\theta_{l})}_{\text{unlabeled data likelihood}}+\underbrace{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\log\sum_{l=1}^{L}\pi_{l}p(\mathbf{f}_{i}|\theta_{l})P(c_{i}|l)}_{\text{labeled data likelihood}}$ (3)
where $\Theta=\{\pi_{1},\ldots,\pi_{L},\theta_{1},\ldots,\theta_{L}\}$ is the set of model parameters with $\sum_{l=1}^{L}\pi_{l}=1$, $\pi_{l}$ represents the mixing coefficient of the $l$-th Gaussian component, $p(\mathbf{f}_{i}|\theta_{l})$ is the probability density function of the $l$-th Gaussian component with parameters $\theta_{l}=\{\bm{\mu}_{l},\Sigma_{l}\}$, and $P(c_{i}|l)$ represents the conditional probability that a feature belongs to class $c_{i}$ given Gaussian component $l$.
The EM algorithm updates parameters as follows:
$\gamma_{il}^{(t)}=\frac{\pi_{l}^{(t)}p(\mathbf{f}_{i}|\theta_{l}^{(t)})}{\sum_{j=1}^{L}\pi_{j}^{(t)}p(\mathbf{f}_{i}|\theta_{j}^{(t)})}\quad\forall\mathbf{f}_{i}\in\mathcal{F}_{u}$ (4)
$\gamma_{il|c_{i}}^{(t)}=\frac{\pi_{l}^{(t)}p(\mathbf{f}_{i}|\theta_{l}^{(t)})P(c_{i}|l)^{(t)}}{\sum_{j=1}^{L}\pi_{j}^{(t)}p(\mathbf{f}_{i}|\theta_{j}^{(t)})P(c_{i}|j)^{(t)}}\quad\forall\mathbf{f}_{i}\in\mathcal{F}_{l}$ (5)
$\bm{\mu}_{l}^{(t+1)}=\frac{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\mathbf{f}_{i}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\mathbf{f}_{i}\gamma_{il}^{(t)}}{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\gamma_{il}^{(t)}}$ (6)
$\Sigma_{l}^{(t+1)}=\frac{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}M_{il}^{(t)}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}M_{il}^{(t)}\gamma_{il}^{(t)}}{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\gamma_{il}^{(t)}}$ (7)
where $M_{il}^{(t)}=(\mathbf{f}_{i}-\bm{\mu}_{l}^{(t)})(\mathbf{f}_{i}-\bm{\mu}_{l}^{(t)})^{\top}$.
$\pi_{l}^{(t+1)}=\frac{1}{N}\left(\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\gamma_{il|c_{i}}^{(t)}+\sum_{\mathbf{f}_{i}\in\mathcal{F}_{u}}\gamma_{il}^{(t)}\right)$ (8)
$P(k|l)^{(t+1)}=\frac{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l},c_{i}=k}\gamma_{il|c_{i}}^{(t)}}{\sum_{\mathbf{f}_{i}\in\mathcal{F}_{l}}\gamma_{il|c_{i}}^{(t)}}$ (9)
3.3 Pseudo-labeling Mechanism
To effectively leverage information from the unlabeled dataset $\mathcal{X}_{u}$, we introduce a pseudo-labeling mechanism based on the component-class association probabilities. Unlike traditional iterative methods, our pseudo-label generation is performed in a single calculation after the initial SGMM training reaches convergence, reducing computational cost and mitigating error accumulation risks.
For unlabeled features $\mathbf{f}_{i}\in\mathcal{F}_{u}$, we compute class probabilities through Gaussian component responsibilities:
$p_{i}^{(k)}=\sum_{l=1}^{L}P(k|l)^{(t)}\gamma_{il}^{(t)},\quad\xi_{i}=\max_{1\leq k\leq K}p_{i}^{(k)}$ (10)
where $P(k|l)$ represents the class probability conditioned on component $l$, and $\gamma_{il}^{(t)}$ is the responsibility from Equation 4.
We construct class-specific candidate sets with confidence thresholding sorted by $\xi_{i}$ descending:
$\mathcal{C}_{k}=\left\{\mathbf{f}_{i}\mid\arg\max_{k}p_{i}^{(k)}=k,\ \xi_{i}>\tau,\quad\tau\in(0,1)\right\}$ (11)
To prevent class imbalance, we adopt proportional sampling:
$n^{\text{final}}=\min_{1\leq k\leq K}\lfloor\alpha|\mathcal{C}_{k}|\rfloor\quad(\alpha\in(0,1))$ (12)
The pseudo-labeled set is then constructed as:
$\mathcal{D}_{p}=\{(\mathbf{f}_{i},k)\mid k=1,\ldots,K,\ \mathbf{f}_{i}\in\mathcal{C}_{k}[1:n^{\text{final}}]\}$ (13)
Finally, combine the pseudo-labeled data $\mathcal{D}_{p}$ with the labeled features $\mathcal{F}_{l}$ and use the EM algorithm for iterative. We use Kmeans++ to initialize the parameters first, during the training process, log-likelihood value calculated using Equation 3 increases steadily. We show the procedure of our method in Algorithm 1.
Algorithm 1 SemiOccam Network
0:  Labeled data $\mathcal{X}_{l}$, Unlabeled data $\mathcal{X}_{u}$, Pre-trained ViT, SGMM components $L$, Confidence threshold $\tau$, Sampling ratio $\alpha$
0:  Trained SGMM parameters $\Theta$
1:  Extract features $\mathcal{F}_{l}\leftarrow\{\text{ViT}(\mathbf{x}_{i})|(\mathbf{x}_{i},c_{i})\in\mathcal{X}_{l}\}$ and $\mathcal{F}_{u}\leftarrow\{\text{ViT}(\mathbf{x}_{i})|\mathbf{x}_{i}\in\mathcal{X}_{u}\}$
2:  PCA $\mathcal{F}_{l}$ and $\mathcal{F}_{u}$ into $\mathbb{R}^{d}$
3:  $L$ centers cluster $\mathcal{F}_{l}\cup\mathcal{F}_{u}$ with K-means++
4:  Initialize $\{\pi_{l},\bm{\mu}_{l},\Sigma_{l}\}$ from cluster results
5:  for $t=1$ to $T_{1}$ do
6:     EM algorithm: Initial iterative via (4-9)
7:  end for
8:  Construct $\mathcal{D}_{p}$ using (11-13) with $\tau$ and $\alpha$
9:  Augment labeled set: $\mathcal{F}_{l}\leftarrow\mathcal{F}_{l}\cup\mathcal{D}_{p}$
10:  for $t=1$ to $T_{2}$ do
11:     EM algorithm: Final iterative via (4-9)
12:  end for
13:  return  $\Theta$
4 Experiments
4.1 Standard SGMM Performance Analysis
The performance of the SGMM is influenced by several factors. This section will delve into the impact of the number of labels, feature dimensions, and the number of Gaussian components on the model’s behavior.
SGMM’s core of its labeling efficiency lies in using unlabeled data to assist the model in learning the data distribution. Theoretically, as the number of labeled samples increases, the model can more accurately estimate the conditional probability of categories $P(k|l)$, thereby improving discrimination ability. When the labeled samples reach a certain scale, the marginal benefit of the additional information provided by unlabeled data diminishes, and the performance improvement curve tends to flatten. Therefore, SGMM’s advantage is that it can quickly improve performance with very few labels and demonstrate high efficiency in applications with limited labeling resources.
Figure 3: Relationship between number of labels/PCA dimensions/Gaussian components and performance on the MNIST dataset. The first figure shows a steady improvement in accuracy with an increase in labels. The second figure shows the trend of SGMM accuracy with PCA dimensions, while the right figure shows the trend of accuracy with the number of Gaussian submodels.
When dealing with high-dimensional data, the feature dimension is crucial to the performance and computational efficiency of SGMM. High-dimensional feature spaces can lead to the “curse of dimensionality”, increasing the complexity of model training and the risk of overfitting, while also consuming excessive computational resources. Reducing feature dimensions through dimensionality reduction techniques such as PCA can retain discriminative feature information while removing redundant dimensions, alleviating these problems to some extent. However, excessive dimensionality reduction can lead to the loss of key feature information. Therefore, the feature dimension should be reduced as much as possible while ensuring model performance.
Gaussian Mixture Models use multiple Gaussian components to fit complex data distributions, and the number of components directly affects the model’s complexity and fitting ability. When the number of components is small, the model may not be able to fully capture the fine-grained structure of the data distribution, leading to underfitting. When the number of components is too large, the model complexity increases, which can easily lead to overfitting of the training data, reducing generalization ability and increasing computational costs. Therefore, the optimal number of Gaussian components should match the intrinsic class structure and complexity of the dataset.
In summary, labeling efficiency reflects the core advantage of semi-supervised learning. Feature dimension efficiency emphasizes the role of dimensionality reduction in improving model efficiency and generalization ability. The number of Gaussian components reflects the trade-off between model complexity and data fitting ability. Understanding the impact of these factors on SGMM performance helps to adjust model parameters according to specific tasks in practical applications.
4.2 More Powerful Feature Extractors
To further analyze the performance of the SGMM, we experimented on the CIFAR-10 dataset using different feature extractors: PCA, Deep Neural Networks (ResNet101, DLA169), and ViT-Base/Large backbone DINO pre-trained models. Figure 4 shows that with DNNs and ViTs, increasing the number of Gaussian components leads to a decrease in classification accuracy, which contrasts with the performance trend observed when using PCA.
Figure 4: Performance trends. The left figure shows the accuracy trend of SGMM on the CIFAR-10 dataset processed with PCA, while the right figure shows the accuracy trend with the number of Gaussian submodels when using DNNs and ViTs.
Next, we performed t-SNE visualization on the features extracted by the three methods. In Figure 5, we present the t-SNE visualization results of feature representations obtained using different feature extraction methods on the CIFAR-10 dataset. It can be clearly seen from the figure that the feature points extracted by PCA exhibit a highly mixed state in the t-SNE space, and it is difficult to distinguish between samples from different categories. This indicates that the feature representation after PCA dimensionality reduction lacks sufficient discriminative power. In contrast, the feature points extracted using deep neural networks exhibit a certain clustering trend in the t-SNE space, but there is still some overlap between the boundaries of different categories. The feature points extracted using Visual Transformer form a relatively clear cluster structure in the t-SNE space, which indicates that the feature representation extracted by ViT has stronger separability and can better capture the intrinsic structure of the data, providing a more favorable feature basis for subsequent classification tasks.
PCA has limited feature discrimination ability, so it needs more Gaussian sub-models to model complex intra-class variations. At this time, increasing the number of Gaussian components helps to approximate the true distribution; while the features extracted by DNN/ViT are close to a single Gaussian for each category, increasing the number of components will cause the model complexity to exceed the actual needs. These analyses are consistent with the experimental results in Figure 4.
Figure 5: t-SNE Visualization. t-SNE visualization results of different feature extraction methods on the CIFAR-10 dataset.
We analyzed high-confidence samples from two Gaussian components of the same bird category. Observing the original images in Figure 6, we found that these two Gaussian components exhibit different feature preferences: one component tends to give high confidence to samples of ostriches (long-necked birds), while the other component prefers plump, neckless birds.
Figure 6: Comparison raw images. Visualization of high-confidence samples from two Gaussian components of the bird category on the CIFAR-10 dataset.
Through experimental results and visualization analysis, we found that DNN and ViT can better capture the distribution characteristics of data compared to the traditional PCA dimensionality reduction method, thereby improving the classification performance of SGMM.
4.3 Experiment Setup
Datasets.
We report results on the Cifar-10/100 and CleanSTL-10 datasets. CIFAR-10 contains 10 classes with 6,000 32x32 color images per class, totaling 60,000 images, of which 50,000 are for training and 10,000 for testing. CIFAR-100 is similar to CIFAR-10 but contains 100 classes with 600 images per class, of which 500 are for training and 100 for testing. STL-10 is an image classification dataset designed for self-supervised and semi-supervised learning, containing 10 classes with 5,000 labeled training samples and 100,000 unlabeled images, each at a resolution of 96×96 pixels. We created CleanSTL-10 by carefully removing duplicates and overlaps between training and test sets, as detailed in Appendix A. After this deduplication, CleanSTL-10 contains 90,455 training images and 8,000 testing images, ensuring more reliable and fair evaluation.
Implementation details.
We use the DINO pre-trained model based on the ViT-large backbone to extract feature representations. On the Cifar-10/100 datasets, the values of the Gaussian components are set to 10/100 respectively, and the PCA dimension is set to 60. In the STL-10 dataset, since its training set contains more than 10 categories, we set the number of Gaussian components to 15 to obtain better robustness, the PCA dimension is set to 45, and the convergence threshold (Tol) is set to 1. The PCA dimension is set to use accuracy as an evaluation metric to evaluate the classification performance of different methods. All experiments are performed on a Tesla T4 GPU with 15GB memory.
4.4 Comparison with State-of-the-art Methods
Results on CIFAR-10.
We show the benchmark results on CIFAR-10 in Table 1. Our method achieves competitive performance compared with state-of-the-art methods, including FixMatch, UDA, and ReMixMatch. It can be clearly seen that our method consistently outperforms other methods under all settings.
Table 1: Benchmark results on CIFAR-10.
Algorithms Error Rate (40 labels) Error Rate (250 labels) Error Rate (4000 labels)
Dash (2021) 9.29 $\pm$ 3.28 5.16 $\pm$ 0.23 4.36 $\pm$ 0.11
MPL (2021) 6.62 $\pm$ 0.91 5.76 $\pm$ 0.24 4.36 $\pm$ 0.11
FlexMatch (2021) 4.97 $\pm$ 0.06 4.98 $\pm$ 0.09 4.19 $\pm$ 0.01
CoMatch (2021) 6.51 $\pm$ 1.18 5.35 $\pm$ 0.14 4.27 $\pm$ 0.12
SimMatch (2022) 5.38 $\pm$ 0.01 5.36 $\pm$ 0.08 4.41 $\pm$ 0.07
AdaMatch (2022) 5.09 $\pm$ 0.21 5.13 $\pm$ 0.05 4.36 $\pm$ 0.05
FreeMatch (2023) 4.90 $\pm$ 0.12 4.88 $\pm$ 0.09 4.16 $\pm$ 0.06
SoftMatch (2023) 5.11 $\pm$ 0.14 4.96 $\pm$ 0.09 4.27 $\pm$ 0.05
SequenceMatch (2024a) 4.80 $\pm$ 0.01 4.75 $\pm$ 0.05 4.15 $\pm$ 0.01
EPASS (2024b) 5.55 $\pm$ 0.21 5.31 $\pm$ 0.13 4.23 $\pm$ 0.05
SemiOccam(Ours) 3.51 $\pm$ 0.12 3.47 $\pm$ 0.17 3.45 $\pm$ 0.16
Results on CleanSTL-10.
In Table 2, we present the performance on CleanSTL-10. Our method achieves state-of-the-art performance, outperforming other methods by a large margin. Our method achieves the best performance in all three settings, demonstrating the effectiveness of our proposed method. Compared to SequenceMatch in 2024a, our SemiOccam achieves +10.88%, +8.35%, and +1.41% on 40-label, 250-label, and 1000-label settings, respectively.
Table 2: Benchmark results on CleanSTL-10.
Algorithms Error Rate (40 labels) Error Rate (250 labels) Error Rate (1000 labels)
Dash (2021) 42.00 $\pm$ 4.94 10.50 $\pm$ 1.37 6.30 $\pm$ 0.49
MPL (2021) 35.97 $\pm$ 4.14 9.90 $\pm$ 0.96 6.66 $\pm$ 0.00
FlexMatch (2021) 29.12 $\pm$ 5.04 9.85 $\pm$ 1.35 6.08 $\pm$ 0.34
CoMatch (2021) 13.74 $\pm$ 4.20 7.63 $\pm$ 0.94 5.71 $\pm$ 0.08
SimMatch (2022) 16.98 $\pm$ 4.24 8.27 $\pm$ 0.40 5.74 $\pm$ 0.31
AdaMatch (2022) 19.95 $\pm$ 5.17 8.59 $\pm$ 0.43 6.01 $\pm$ 0.02
FreeMatch (2023) 28.50 $\pm$ 5.41 9.29 $\pm$ 1.24 5.81 $\pm$ 0.32
SoftMatch (2023) 22.23 $\pm$ 3.82 9.18 $\pm$ 0.63 5.79 $\pm$ 0.15
SequenceMatch (2024a) 15.45 $\pm$ 1.40 12.78 $\pm$ 0.76 5.56 $\pm$ 0.35
EPASS (2024b) 9.15 $\pm$ 3.25 6.27 $\pm$ 0.03 5.40 $\pm$ 0.12
SemiOccam(Ours) 4.57 $\pm$ 0.24 4.43 $\pm$ 0.08 4.15 $\pm$ 0.07
Results on CIFAR-100.
In Table 3, we compare the performance of SemiOccam with state-of-the-art methods on CIFAR-100 dataset. Our method outperforms other methods in the 400 labels and 2,500 labels setting, and achieves comparable 10,000 labels settings.
Table 3: Benchmark results on CIFAR-100.
Algorithms Error Rate (400 labels) Error Rate (2500 labels) Error Rate (10000 labels)
Dash (2021) 44.82 $\pm$ 0.96 27.15 $\pm$ 0.22 21.88 $\pm$ 0.07
MPL (2021) 46.26 $\pm$ 1.84 27.71 $\pm$ 0.19 21.74 $\pm$ 0.09
FlexMatch (2021) 39.94 $\pm$ 1.62 26.49 $\pm$ 0.20 21.90 $\pm$ 0.15
CoMatch (2021) 53.41 $\pm$ 2.36 29.78 $\pm$ 0.11 22.11 $\pm$ 0.22
SimMatch (2022) 39.32 $\pm$ 0.72 26.21 $\pm$ 0.37 21.50 $\pm$ 0.11
AdaMatch (2022) 38.08 $\pm$ 1.35 26.66 $\pm$ 0.33 21.99 $\pm$ 0.15
FreeMatch (2023) 39.52 $\pm$ 0.01 26.22 $\pm$ 0.08 21.81 $\pm$ 0.17
SoftMatch (2023) 37.60 $\pm$ 0.24 26.39 $\pm$ 0.38 21.86 $\pm$ 0.16
SequenceMatch (2024a) 37.86 $\pm$ 1.07 25.99 $\pm$ 0.22 20.10 $\pm$ 0.04
EPASS (2024b) 38.88 $\pm$ 0.24 25.68 $\pm$ 0.33 21.32 $\pm$ 0.14
SemiOccam(Ours) 26.59 $\pm$ 1.02 22.19 $\pm$ 0.81 21.21 $\pm$ 0.26
4.5 Training Cost Analysis
In the field of semi-supervised learning, besides model performance, training overhead is a key metric for evaluating the practicality and environmental impact of algorithms. This section will quantitatively compare the training overhead of existing benchmarks with our proposed method.
TorchSSL from Zhang et al. is a mainstream SSL evaluation protocols, typically employ training deep neural networks from scratch, leading to significant computational burdens. For example, training the FixMatch algorithm on a single NVIDIA V100 GPU requires “300 hours $\times$ 3 settings $\times$ 3 seeds” for the CIFAR-100 dataset; “110 hours $\times$ 3 settings $\times$ 3 seeds” for CIFAR-10; and “225 hours $\times$ 3 settings $\times$ 3 seeds” for STL-10. The total training time amounts to 5715 GPU hours (approximately 238 GPU days). This high computational demand poses a significant experimental cost barrier for academia and resource-constrained research institutions.
USB from Wang et al. is another current mainstream SSL evaluation protocol. It reports that training the FixMatch algorithm on a single NVIDIA V100 GPU incurs an overhead of “11 hours $\times$ 2 settings $\times$ 3 seeds” for the CIFAR-100 dataset, and “18 hours $\times$ 2 settings $\times$ 3 seeds” for the STL-10 dataset. Across these two datasets, a total of 174 GPU hours are required.
Our proposed method utilizes a pre-trained ViT as a feature extractor. After this feature extraction step is completed, the core SGMM classifier training process can be completed in less than ten minutes on a free Tesla T4 GPU 15GB on the Google Colab platform, demonstrating high efficiency and state-of-the-art performance. This minute-level training time gives our method a significant competitive advantage in resource-constrained practical deployments and rapid iteration research scenarios.
4.6 Ablation Study
Comparing Semi-supervised Learning Methods.
In order to test the effectiveness of the SGMM component, we compared it with the original classification head of ViT. We trained on three datasets, error rates shown in Table 4, under scenarios with extremely few labels, the performance of SGMM is significantly superior to that of the Softmax layer.
Effect of Pseudo-labeling.
We conducted an ablation study to evaluate the impact of pseudo-labeling on the performance of SemiOccam. We compared the performance of SemiOccam with and without pseudo-labeling on three datasets, error rates shown in Table 4.
Table 4: Ablation study with three datasets.
Dataset CIFAR-10 CIFAR-100 CleanSTL-10
# Label 40 250 4000 400 2500 10000 40 250 1000
w Softmax 51.05 $\pm$ 16.69 5.40 $\pm$ 1.70 3.75 $\pm$ 0.21 57.32 $\pm$ 10.91 43.01 $\pm$ 5.89 33.70 $\pm$ 2.71 16.29 $\pm$ 2.85 4.88 $\pm$ 1.50 4.38 $\pm$ 0.40
w/o P-L 3.73 $\pm$ 0.21 3.70 $\pm$ 0.23 3.63 $\pm$ 0.19 30.56 $\pm$ 0.80 23.21 $\pm$ 0.66 21.56 $\pm$ 0.43 4.91 $\pm$ 0.12 4.81 $\pm$ 0.12 4.58 $\pm$ 0.10
SemiOccam 3.51 $\pm$ 0.12 3.47 $\pm$ 0.17 3.45 $\pm$ 0.16 26.59 $\pm$ 1.02 22.19 $\pm$ 0.81 21.21 $\pm$ 0.26 4.44 $\pm$ 0.04 4.43 $\pm$ 0.08 4.15 $\pm$ 0.07
5 Conclusions
We introduce high-performance feature extractors into the classical yet often underappreciated MoE classifier for image recognition tasks. Our architecture is simple and training is highly efficient, making it particularly well-suited for scenarios with limited labeled data. The proposed unified modeling strategy performs surprisingly well when trained with very few labeled samples and large-scale unlabeled data. In our study, we also identify a long-overlooked data leakage issue in the STL-10 dataset and release the deduplicated CleanSTL-10 dataset to ensure fair and reliable evaluation in semi-supervised learning. Extensive experiments show that SemiOccam achieves new performance breakthroughs on benchmarks such as CleanSTL-10 and CIFAR-10, maintaining over 96% classification accuracy even with only 4 labeled samples per class in under ten minutes on a single Tesla T4 GPU, compared to the hundreds of GPU hours required by existing methods.
Future development of this technology can proceed along several dimensions. First, constructing dynamic component adjustment mechanism based on a Bayesian framework to achieve parameter adaptivity. Secondly, enhancing the model’s discriminative performance when facing complex tasks such as fine-grained classification. Most interesting, exploring its application in other fields, such as electroencephalogram signal analysis and medical image processing, thereby broadening its application scope and verifying its transferability.
We believe these improvements will further unlock the technological benefits brought about by the fusion of probabilistic graphical models and deep learning. Crucially, given our method’s advantages in short training times and high efficiency, we are particularly keen to see its widespread adoption in edge deployment scenarios and applications requiring rapid iteration.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62076078 and in part by the Chinese Association for Artificial Intelligence (CAAI)-Huawei MindSpore Open Fund under Grant CAAIXSJLJJ-2020-033A.
APPENDIX
Appendix A STL-10 Data Clean
Duplicate samples in the STL-10 dataset can potentially bias model evaluation. To address this, we deduplicated the dataset. Specifically, we used a script to compute image hashes for STL-10 dataset, then removed 7,545 samples from the train set that were found to be duplicates of samples in the test set. We release the deduplicated CleanSTL-10 dataset to facilitate fair and reliable research. The pseudo-code for this deduplication process is provided in Algorithm A.1.
Algorithm A.1 Deduplication of Dataset
0:  Main dataset, test dataset, batch size
0:  Deduplicated dataset, CSV file with duplicate information
  Build test set hash dictionary: Compute hash for each test image
2:  Initialize valid_indices and duplicates as empty lists
  for each batch in main dataset do
4:     for each image in batch do
        Compute hash of the image
6:        if hash exists in test set hash dictionary then
           Record duplicate sample details
8:        else
           Add current index to valid_indices
10:        end if
     end for
12:  end for
  if duplicates is not empty then
14:     Save duplicate details to CSV file
  end if
Following processing, the training set consists of 5,000 labeled samples and 90,455 unsupervised samples. This effectively addresses the problem of data leakage, thereby guaranteeing the credibility of the experimental findings.
We have compiled a table of the indices for training and test samples with duplicate hash values, which you can verify in our project repository.
Figure B.1: Heatmap comparison. From left to right: raw image, DLA169, DINO. Note the small penguin on the right in the third row.
Appendix B Comparing Heatmaps of DNN and ViT
We compare the heatmaps of DNN and ViT on several randomly selected images from the internet, each containing two animals simultaneously. The heatmaps are generated using the Grad-CAM method, which visualizes the importance of each pixel in the image for the classification result. As shown in Figure B.1, the heatmaps of DNN and ViT differ significantly. While DNN focuses primarily on one animal, ViT attends to both animals in the image, demonstrating its ability to capture global context more effectively. This comparison highlights the differences in how these models prioritize regions of the image during classification.
Appendix C PCA Explained Variance
In our experiments, the hyperparameter of PCA dimensions for different datasets is determined by the PCA explained variance, which represents the percentage of variance in the data that can be explained by the feature vectors after PCA dimensionality reduction. We typically choose dimensions that can explain over 60% of the variance.
Figure C.1: PCA explained variance. The percentage of variance explained by each PCA dimension on the CIFAR-10 dataset.
Appendix D ViT Classification Head for Semi-Supervised Learning
The following is the mathematical implementation of the original ViT classification head for semi-supervised learning, which has been used in our ablation study in section 4.6.
In section 3.1, we define the feature vector $\mathbf{f}_{i}\in\mathbb{R}^{d}$. The standard ViT classification head includes a linear transformation $\mathbf{s}_{i}=\mathbf{W}_{h}\mathbf{f}_{i}+\mathbf{b}_{h}$, where $\mathbf{W}_{h}\in\mathbb{R}^{K\times d}$ and $\mathbf{b}_{h}\in\mathbb{R}^{K}$ are learnable parameters, and $K$ is the number of categories; followed by softmax:
$P(c=k|\mathbf{f}_{i})=\frac{\exp(s_{i,k})}{\sum_{j=1}^{K}\exp(s_{i,j})},\quad\forall k\in\{1,\ldots,K\}$
For labeled samples $\mathcal{F}_{l}=\{(\mathbf{f}_{i},c_{i})\}$, we optimize the standard cross-entropy loss:
$\mathcal{L}_{\text{sup}}=-\frac{1}{|\mathcal{F}_{l}|}\sum_{(\mathbf{f}_{i},c_{i})\in\mathcal{F}_{l}}\log P(c=c_{i}|\mathbf{f}_{i})$
Unlabeled data $\mathcal{F}_{u}=\{\mathbf{f}_{j}\}$ adopts a confidence-based pseudo-labeling strategy, where the pseudo-label is $\hat{c}_{j}=\arg\max_{k}P(c=k|\mathbf{f}_{j})$, and the confidence score is $\xi_{j}=\max_{k}P(c=k|\mathbf{f}_{j})$. The unsupervised loss is calculated as follows:
$\mathcal{L}_{\text{unsup}}=-\frac{1}{|\mathcal{F}_{u}|}\sum_{\mathbf{f}_{j}\in\mathcal{F}_{u}}\mathbb{I}(\xi_{j}>\tau)\log P(c=\hat{c}_{j}|\mathbf{f}_{j})$
The total loss combines the supervised and unsupervised objectives, with the weight $\lambda(t)$ changing over time, where $T_{\text{ramp}}$ corresponds to the training period:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{sup}}+\underbrace{\lambda_{\text{max}}\cdot\min\left(1,\frac{t}{T_{\text{ramp}}}\right)}_{\lambda(t)}\mathcal{L}_{\text{unsup}}$